# Cosmology with voids

### v3

@@ 1 Introduction
Cosmic voids fill the universe. The vast, empty voids span tens of megaparsecs and contain few, if any, bright galaxies. Increasingly ambitious panoramic redshift surveys reveal the cosmic web with filaments, sheets and clusters of galaxies delineating a sea of voids (e.g., Geller & Huchra, 1989; Strauss et al., 2002; Shectman et al., 1996; Colless et al., 2003; Ahn et al., 2012; Liske et al., 2015; Drinkwater et al., 2018; Sohn et al., 2023; Hahn et al., 2023).
@@
Despite their low mass density, voids play a key role in shaping dense regions of the universe (Bertschinger, 1985; White et al., 1987; Regos & Geller, 1991; Ceccarelli et al., 2013). They grow relative to the cosmic expansion. As they grow, they push matter onto shells at their outer boundaries. Their growth also feeds into galaxy clusters and superclusters. Like the massive, collapsed structures, voids reflect the “background cosmology”, the mean density of matter and dark energy, and the overall expansion rate. They also contain the signature of the primordial spectrum of density fluctuations, the seeds of the large-scale structure observed today.
@@
Early theoretical work considered void evolution (Alcock & Paczynski, 1979; Aarseth & Saslaw, 1982; Hoffman & Shaham, 1982; Bertschinger, 1985; Melott, 1987; White et al., 1987; Regos & Geller, 1991; Kauffmann & Fairall, 1991; Blumenthal et al., 1992), while observational investigations (El-Ad & Piran, 1997; Kauffmann & Fairall, 1991; Pan et al., 2012; Sutter et al., 2012; Micheletti et al., 2014; Szapudi et al., 2015; Finelli et al., 2016; Mackenzie et al., 2017; Kovács et al., 2022; Owusu et al., 2023; Douglass et al., 2023) yielded void catalogs. More recent theoretical studies explored the void size distribution (Kamionkowski et al., 2009; Sahlén et al., 2016; Sahlén & Silk, 2018; Correa et al., 2021; Contarini et al., 2022; Verza et al., 2024), ellipticity distribution (Lavaux & Wandelt, 2010), density and velocity profiles (Hamaus et al., 2014a, b; Massara & Sheth, 2018; Stopyra et al., 2021; Voivodic et al., 2020; Voivodic & Barreira, 2021; Verza et al., 2022), and redshift-space distortions (Hamaus et al., 2015; Chuang et al., 2017; Cai et al., 2016; Nadathur & Percival, 2019), including the Alcock-Paczynski (AP) test (Hamaus et al., 2022; Radinović et al., 2023) based on the expected statistical isotropy of large-scale structure. Other work emphasized the connection between void statistics and cosmological models (Spolyar et al., 2013; Clampitt et al., 2013; Cai et al., 2015; Massara et al., 2015; Zivick et al., 2015; Voivodic et al., 2017; Kreisch et al., 2019; Perico et al., 2019; Verza et al., 2019; Contarini et al., 2021; Bayer et al., 2021; Kreisch et al., 2022; Pelliciari et al., 2023; Verza et al., 2023; Wilson & Bean, 2023; Vielzeuf et al., 2023), and considered void properties in simulations (Nadathur & Hotchkiss, 2015a, b; Nadathur et al., 2017; Pollina et al., 2016; Schuster et al., 2023, 2024). Analyses of void data, based on redshift-space distortions and the AP test (Mao et al., 2017; Hamaus et al., 2016, 2017; Hawken et al., 2017; Achitouv et al., 2017; Achitouv, 2019; Nadathur et al., 2019b, 2020c; Aubert et al., 2022; Hamaus et al., 2020; Hawken et al., 2020), along with the void size function (Contarini et al., 2023), demonstrate the importance of voids as cosmological indicators.
@@
Although linear perturbation theory for the growth of cosmic structure (Peebles, 1980) may describe the evolution of shallow voids on large scales, numerical simulations are required to connect theory and properties of voids as they emerge in the cosmic web. A challenge is that simulation volumes must be large with length scales of $\sim$1 Gpc to track “typical” superclusters and large voids. Most current hydrodynamical/$n$-body simulations incorporate cosmological parameters that best reflect the observations (Springel et al., 2005; Angulo et al., 2012; Klypin et al., 2011; McCarthy et al., 2017; Lee et al., 2021; Kugel et al., 2023). While not used for fitting cosmological parameters, these simulations provide model tests and may reveal tension between observations and theory.
@@
Here we revisit the cosmological information contained in large voids. We confirm that void properties, including their size, interior mass density, and velocity flows, are sensitive to parameters of the background cosmology (Blumenthal et al., 1992). Void properties alone may provide estimates of these parameters, drawn from a range of cosmological models with baryons, dark matter and dark energy. By connecting the properties of voids to the broader universe, we test whether their dynamics are well-described by standard cosmology. Voids may eventually provide new insight into dark energy from rarefied environments where its dynamical effects are presumably greater than in the universe at large.
@@
We follow two complementary approaches to understand voids in the cosmological context. Following Blumenthal et al. (1992), we examine the evolution of idealized spherical voids numerically. This approach lays the foundation for relating the void depths and outflows to the underlying cosmology. We refine the cosmological constraints provided by the analytic analysis by generating a large suite of $n$-body models that treat the growth of large-scale structure in the quasilinear regime (Zel’dovich, 1970; White, 2014; Nadathur et al., 2019a; Stopyra et al., 2021; Schuster et al., 2023). This application of the adhesion approximation (Weinberg & Gunn, 1990) enables the realization of snapshots of large-scale structure in patches of the universe along with the construction of lightcones that mimic large redshift surveys. We demonstrate that in theory, when the entire phase space is available, voids place tight constraints on the underlying cosmology. Any significant difference between these constraints from voids and other cosmological measurements may provide insight into the dynamical effects of dark energy.
@@
We focus on the dynamics of matter flow within voids where the ratio of dark matter density to dark energy density is generally low. Other work has considered the interface between (linear) voids and (nonlinear) neighboring high-density regions, including void shapes and void-galaxy cross correlations (Alcock & Paczynski, 1979; Lee & Park, 2009; Hamaus et al., 2015; Cai et al., 2016; Nadathur et al., 2020a; Woodfinden et al., 2022). These analyses often depend on one or a few large-scale $n$-body simulations of the standard cosmology to compare theory and observation. Here we explore a wide range of models, sampling over 100 realizations of the cosmic density and velocity fields with $10^{7}$$10^{9}$ particles to assess sensitivity to a broad range of cosmological parameters.
@@
Section 2 describes the suite of cosmological models we use as a test bed. In section 3, we examine idealized voids to assess the role of matter and dark energy in void evolution. We use linear theory to interpret the results and we extract cosmological parameters from simple void properties (section 4). In section 5, we use a fast $n$-body algorithm, the adhesion approximation (Weinberg & Gunn, 1990), to sample void properties in more realistic scenarios. We assess these simulated voids as probes of the parameters of the background cosmology (section 5.2). In section 6 we consider voids in currently observable redshift space and highlight important differences between the properties of voids in real and redshift space. We conclude in section 7.
@@ 2 Background cosmologies
To explore the cosmological information content of voids, we work with the standard parameters of the Friedmann-Lemâitre-Robertson-Walker (FLRW) framework. The overall expansion of the universe depends on a matter density parameter $\Omega_{\text{M}}$, a parameter $\Omega_{\Lambda}$ associated with dark energy, and the Hubble parameter $H_{0}$. All three parameters refer to values at the present epoch (time $t=t_{\text{now}}$). The mass density $\Omega_{\text{M}}$ includes both baryons and cold dark matter; we approximate the effects of dark energy (Albrecht et al., 2006) with $\Omega_{\Lambda}$ as defined by a cosmological constant $\Lambda$. The set of parameters $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$$H_{0}$ in units of 100 km/s/Mpc — define the “background cosmology”. The cosmic web emerges as a pattern within this background cosmology.
@@
${\dot{a}}=H_{0}\sqrt{a^{-1}\Omega_{\text{M}}+\Omega_{\text{K}}+a^{2}\Omega_{\Lambda}},$ (1)
${\ddot{a}}={H_{0}^{2}}\left[-a^{-2}\Omega_{\text{M}}/2+a\Omega_{\Lambda}\right)$ (2)
@@
$\Omega_{\text{M}}=\frac{8\pi G}{3H_{0}^{2}}\bar{\rho}_{0},$ (3)
$\Omega_{\Lambda}=\frac{c^{2}}{3H_{0}^{2}}\Lambda,$ (4)
@@
${\Omega_{\text{M}}(z)}=(1+z)^{3}\Omega_{\text{M}}H_{0}^{2}/{H^{2}(z)}$ (7)
${\Omega_{\Lambda}(z)}=\Omega_{\Lambda}H_{0}^{2}/{H^{2}(z)}$ (8)
${H(z)}\equiv{\dot{a}}/{a}\hskip 155.38005pt$
$\ \ \ \ =H_{0}\sqrt{(1+z)^{3}\Omega_{\text{M}}+(1+z)^{2}\Omega_{\text{K}}+\Omega_{\Lambda}}.$ (9)
@@ 2.1 Density fluctuations and the matter power spectrum
$\delta(\vec{r})=\frac{\rho(\vec{r})}{\overline{\rho}}-1,$ (10)
where $\rho$ is the local matter density and $\vec{r}$ is a proper displacement, a position vector relative to some arbitrary origin. The power is the Fourier transform of the two-point correlation function of this density contrast and it measures the variance of mode amplitudes for each wavevector $\vec{k}$ when averaged within some large volume of the Universe.
In the early expanding universe, self-gravity and the hydrodynamics of baryon-photon interactions mold primordial fluctuations from quantum effects. The gravitational growth of small density fluctuations initially follows linear perturbation theory (Peebles, 1980),
$\delta(\vec{x},t)=\frac{D(t)}{D(t_{i})}\delta(\vec{x},t_{i}),$ (11)
@@
for cosmologies parameterized with a cosmological constant (Heath, 1977; Percival, 2005). Thus, given a primordial density field, we may scale the fluctuation mode amplitudes, and the power spectrum itself, to any epoch we choose. By convention, the primordial power spectrum refers to the epoch just after photons and baryons decouple in the expanding, cooling universe, but before density fluctuation mode amplitudes grow close to unity. This primordial power is typically scaled to the present epoch.
@@
$W_{TH}(k,R)=\frac{3}{(kR)^{2}}\left[\frac{\sin(kR)}{kR}-\cos(kR)\right]$ (14)
is the Fourier transform of a spherical tophat function of radius $R$, normalized so that it yields unity when integrated over space. Equation (13) gives the standard deviation of fluctuation amplitudes when sampled with a tophat window function of radius $R$; by convention, $R=8$ $h^{-1}$pc  is chosen when quantifying the amplitude of fluctuations in a cosmological model. Other choices, for example $R=50$ $h^{-1}$pc, offer insight into the shape of the power spectrum beyond its normalization.
To calculate $P(k)$ for particular model parameters, we use the CAMB software package111github.com/cmbant/CAMB by Lewis et al. (2000a), based on the method of Seljak & Zaldarriaga (1996) and the fast analytical prescriptions of Eisenstein & Hu (1998), as implemented in the nbodykit package222github.com/bccp/nbodykit.
@@ 2.2 Summary of model parameters
[…]
@@
The selected cosmological models are defined with $\Omega_{\text{M}}$, $\Omega_{\Lambda}$, and $\Omega_{\text{b}}$, as in the text, and $\Omega_{\text{K}}$ follows from $1=\Omega_{\text{M}}+\Omega_{\Lambda}+\Omega_{\text{K}}$. The $\sigma_{\text{R}}$ columns refer to the standard deviation of the primordial matter density fluctuations averaged in tophat window functions of radius $R$ (equation (13)) evaluated at 8 $h^{-1}$pc and 50 $h^{-1}$pc. The final column is the age of each model, $t_{\text{now}}$. All parameters are evaluated at this age, including the fluctuation amplitudes.
@@ 3 Idealized voids
An informative starting point for exploring the connection between voids and their background cosmology is an idealized spherical region with a constant density that is, at some early time, slightly less than the average density in the universe. In isolation, this idealized void evolves as if were its own miniature universe with a relatively small matter density and large Hubble parameter (Bertschinger, 1994, 1985). When embedded in a uniform density background universe, we expect the void dynamics to be more complicated as material in the outermost regions of the more rapidly expanding void mixes with the background. Yet the bulk of the idealized void interior evolves as a distinct FLRW world.
We begin this section by considering three approaches to tracking the evolution of idealized voids. The first (section 3.1) comes from linear perturbation theory (Peebles, 1980), valid when the amplitude of the relative density contrast is small ($|\delta|\ll 1$). Because the average density contrast never exceeds unity within a void, linear theory offers a useful approximation to the evolution of the matter density there. The second approach (section 3.2) incorporates the Zel’dovich approximation (Zel’dovich, 1970), a surprisingly robust quasilinear method (White, 2014) for tracking void evolution. Finally, to approach calculating exact solutions, we use the “shell model” described by Blumenthal et al. (1992) to numerically solve the equations of motion for mass elements in an idealized void (section 3.3). These numerical solutions provide an accurate accounting of the mass density even as the outer shells cross into the background universe. Following a brief comparison between these three approaches, we close this section with an assessment how the matter distribution within idealized voids evolve in different background cosmologies.
Throughout, we quantify our results with a void depth parameter (Blumenthal et al., 1992)
$\varepsilon(r)=1-\frac{\left<\rho\right>_{r}}{\bar{\rho}},$ (15)
@@ 3.1 Linear theory
Linear perturbation theory for the growth of cosmic structure (Peebles, 1980) describes the evolution of the mass density when relative fluctuations are small, with $|\delta|\ll 1$ (equation (10)). Here, we provide the predictions of the theory.
@@
$\varepsilon_{\text{lin}}(z)=\frac{D(z)}{D(z_{\text{init}})}\varepsilon(z_{\text{init}}),$ (16)
where redshift $z_{\text{init}}$ is some initial (or reference) redshift, and the denisty growth factors $D$ are defined in equation (12). The theory also predicts flow rates for matter within a void. Including both the Hubble flow and the growing-mode solutions from linear theory, the proper radial velocity relative to the center of a void is (Peebles, 1980)
$v_{r}={H(z)}r+\frac{2fg}{3{\Omega_{\text{M}}(z)}{H(z)}},$ (17)
@@
$=$ $+\frac{{H^{2}(z)}{\Omega_{\text{M}}(z)}r}{2}\varepsilon.$ (20)
The void depth $\varepsilon$ in the last equation refers to a tophat underdensity (equation (15)) or the amplitude of the mean underdensity within comoving radius $x=r/a$. (Equivalent expressions, as in Hamaus et al. 2015, Cai et al. 2015, and Nadathur et al. 2020a, are often written in terms of $\Delta(x)$, the mean density contrast within comoving distance $x$.)
An effective Hubble parameter, ${H_{\text{eff}}}$, describes the radial outflow rate of matter in the interior of an idealized void. The linear theory prediction for ${H_{\text{eff}}}$ in units of the background Hubble expansion rate is (Peebles, 1980; Kaiser, 1987)
$\frac{{H_{\text{eff}}}}{H}$ $=$ $\frac{v_{r}}{r{H(z)}}=1+\frac{2fg}{3{\Omega_{\text{M}}(z)}{H^{2}(z)}r},$ (21)
$=$ $1+\frac{f\varepsilon}{3},$ (22)
which follows from equations (17) and (20). In section 4 we show that this connection between the outflow rate, the growth factor, and void depth leads to a robust estimator of the background cosmology.

### v4

@@ 1 Introduction
Cosmic voids fill the universe. The vast, empty voids span tens of megaparsecs and contain few, if any, bright galaxies. Increasingly ambitious panoramic redshift surveys reveal the cosmic web with filaments, sheets and clusters of galaxies delineating a sea of voids, e.g., [1, 2, 3, 4, 5, 6, 7, 8, 9].
@@
Despite their low mass density, voids play a key role in shaping dense regions of the universe [10, 11, 12, 13]. They grow relative to the cosmic expansion. As they grow, they push matter onto shells at their outer boundaries. Their growth also feeds into galaxy clusters and superclusters. Like the massive, collapsed structures, voids reflect the “background cosmology”, the mean density of matter and dark energy, and the overall expansion rate. They also contain the signature of the primordial spectrum of density fluctuations, the seeds of the large-scale structure observed today.
@@
Early theoretical work considered void evolution [14, 15, 16, 10, 17, 11, 12, 18, 19], while observational investigations [20, 18, 21, 22, 23, 24, 25, 26, 27, 28, 29] yielded void catalogs. More recent theoretical studies explored the void size distribution [30, 31, 32, 33, 34, 35], ellipticity distribution [36], density and velocity profiles [37, 38, 39, 40, 41, 42, 43], and redshift-space distortions [44, 45, 46, 47], including the Alcock-Paczynski (AP) test [48, 49] based on the expected statistical isotropy of large-scale structure. Other work emphasized the connection between void statistics and cosmological models [50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65], and considered void properties in simulations [66, 67, 68, 69, 70, 71]. Analyses of void data, based on redshift-space distortions and the AP test [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82], along with the void size function [83], demonstrate the importance of voids as cosmological indicators.
@@
Although linear perturbation theory for the growth of cosmic structure [84] may describe the evolution of shallow voids on large scales, numerical simulations are required to connect theory and properties of voids as they emerge in the cosmic web. A challenge is that simulation volumes must be large with length scales of $\sim$1 Gpc to track “typical” superclusters and large voids. Most current hydrodynamical/$n$-body simulations incorporate cosmological parameters that best reflect the observations [85, 86, 87, 88, 89, 90]. While not used for fitting cosmological parameters, these simulations provide model tests and may reveal tension between observations and theory.
@@
Here we revisit the cosmological information contained in large voids. We confirm that void properties, including their size, interior mass density, and velocity flows, are sensitive to parameters of the background cosmology [19]. Void properties alone may provide estimates of these parameters, drawn from a range of cosmological models with baryons, dark matter and dark energy. By connecting the properties of voids to the broader universe, we test whether their dynamics are well-described by standard cosmology. Voids may eventually provide new insight into dark energy from rarefied environments where its dynamical effects are presumably greater than in the universe at large.
@@
We follow two complementary approaches to understand voids in the cosmological context. Following [19], we examine the evolution of idealized spherical voids numerically. This approach lays the foundation for relating the void depths and outflows to the underlying cosmology. We refine the cosmological constraints provided by the analytic analysis by generating a large suite of $n$-body models that treat the growth of large-scale structure in the quasilinear regime [91, 92, 93, 40, 70]. This application of the adhesion approximation [94] enables the realization of snapshots of large-scale structure in patches of the universe along with the construction of lightcones that mimic large redshift surveys. We demonstrate that in theory, when the entire phase space is available, voids place tight constraints on the underlying cosmology. Any significant difference between these constraints from voids and other cosmological measurements may provide insight into the dynamical effects of dark energy.
@@
We focus on the dynamics of matter flow within voids where the ratio of dark matter density to dark energy density is generally low. Other work has considered the interface between (linear) voids and (nonlinear) neighboring high-density regions, including void shapes and void-galaxy cross correlations [14, 95, 44, 46, 96, 97]. These analyses often depend on one or a few large-scale $n$-body simulations of the standard cosmology to compare theory and observation. Here we explore a wide range of models, sampling over 100 realizations of the cosmic density and velocity fields with $10^{7}$$10^{9}$ particles to assess sensitivity to a broad range of cosmological parameters.
@@
Section 2 describes the suite of cosmological models we use as a test bed. In section 3, we examine idealized voids to assess the role of matter and dark energy in void evolution. We use linear theory to interpret the results and we extract cosmological parameters from simple void properties (section 4). In section 5, we use a fast $n$-body algorithm, the adhesion approximation [94], to sample void properties in more realistic scenarios. We assess these simulated voids as probes of the parameters of the background cosmology (section 5.2). In section 6 we consider voids in currently observable redshift space and highlight important differences between the properties of voids in real and redshift space. We conclude in section 7.
@@ 2 Background cosmologies
To explore the cosmological information content of voids, we work with the standard parameters of the Friedmann-Lemâitre-Robertson-Walker (FLRW) framework. The overall expansion of the universe depends on a matter density parameter $\Omega_{\text{M}}$, a parameter $\Omega_{\Lambda}$ associated with dark energy, and the Hubble parameter $H_{0}$. All three parameters refer to values at the present epoch (time $t=t_{\text{now}}$). The mass density $\Omega_{\text{M}}$ includes both baryons and cold dark matter; we approximate the effects of dark energy [98] with $\Omega_{\Lambda}$ as defined by a cosmological constant $\Lambda$. The set of parameters $\Omega_{\text{M}}$, $\Omega_{\Lambda}$ and $h$$H_{0}$ in units of 100 km/s/Mpc — define the “background cosmology”. The cosmic web emerges as a pattern within this background cosmology.
@@
${\dot{a}}=H_{0}\sqrt{a^{-1}\Omega_{\text{M}}+\Omega_{\text{K}}+a^{2}\Omega_{\Lambda}},$ (2.1)
${\ddot{a}}={H_{0}^{2}}\left[-a^{-2}\Omega_{\text{M}}/2+a\Omega_{\Lambda}\right)$ (2.2)
@@
$\Omega_{\text{M}}=\frac{8\pi G}{3H_{0}^{2}}\bar{\rho}_{0},$ (2.3)
$\Omega_{\Lambda}=\frac{c^{2}}{3H_{0}^{2}}\Lambda,$ (2.4)
@@
${\Omega_{\text{M}}(z)}=(1+z)^{3}\Omega_{\text{M}}H_{0}^{2}/{H^{2}(z)}$ (2.7)
${\Omega_{\Lambda}(z)}=\Omega_{\Lambda}H_{0}^{2}/{H^{2}(z)}$ (2.8)
${H(z)}\equiv{\dot{a}}/{a}=H_{0}\sqrt{(1+z)^{3}\Omega_{\text{M}}+(1+z)^{2}\Omega_{\text{K}}+\Omega_{\Lambda}}.$ (2.9)
@@ 2.1 Density fluctuations and the matter power spectrum
$\delta(\vec{r})=\frac{\rho(\vec{r})}{\overline{\rho}}-1,$ (2.10)
where $\rho$ is the local matter density and $\vec{r}$ is a position vector relative to some arbitrary origin. The power is the Fourier transform of the two-point correlation function of this density contrast and it measures the variance of mode amplitudes for each wavevector $\vec{k}$ when averaged within some large volume of the Universe.
In the early expanding universe, self-gravity and the hydrodynamics of baryon-photon interactions mold primordial fluctuations from quantum effects. The gravitational growth of small density fluctuations initially follows linear perturbation theory [84],
$\delta(\vec{x},t)=\frac{D(t)}{D(t_{i})}\delta(\vec{x},t_{i}),$ (2.11)
@@
for cosmologies parameterized with a cosmological constant [99, 100]. Thus, given a primordial density field, we may scale the fluctuation mode amplitudes, and the power spectrum itself, to any epoch we choose. By convention, the primordial power spectrum refers to the epoch just after photons and baryons decouple in the expanding, cooling universe, but before density fluctuation mode amplitudes grow close to unity. This primordial power is typically scaled to the present epoch.
@@
$W_{TH}(k,R)=\frac{3}{(kR)^{2}}\left[\frac{\sin(kR)}{kR}-\cos(kR)\right]$ (2.14)
is the Fourier transform of a spherical tophat function of radius $R$, normalized so that it yields unity when integrated over space. Equation (2.13) gives the standard deviation of fluctuation amplitudes when sampled with a tophat window function of radius $R$; by convention, $R=8$ $h^{-1}$pc  is chosen when quantifying the amplitude of fluctuations in a cosmological model. Other choices, for example $R=50$ $h^{-1}$pc, offer insight into the shape of the power spectrum beyond its normalization.
To calculate $P(k)$ for particular model parameters, we use the CAMB software package [^1] The Python version of CAMB is at github.com/cmbant/CAMB. [101], based on the method of [102] and the fast analytical prescriptions of [103], as implemented in the nbodykit package [^2] The nbodykit package is available at github.com/bccp/nbodykit..
@@ 2.2 Summary of model parameters
name $\Omega_{\text{M}}$ $\Omega_{\Lambda}$ $\Omega_{\text{b}}$ $\Omega_{\text{K}}$ $h$ $\sigma_{8}$ $\sigma_{50}$ $t_{\text{now}}$/Gyr
@@
The selected cosmological models are defined with $\Omega_{\text{M}}$, $\Omega_{\Lambda}$, and $\Omega_{\text{b}}$, as in the text, and $\Omega_{\text{K}}$ follows from $1=\Omega_{\text{M}}+\Omega_{\Lambda}+\Omega_{\text{K}}$. The $\sigma_{\text{R}}$ columns refer to the standard deviation of the primordial matter density fluctuations averaged in tophat window functions of radius $R$ (equation (2.13)) evaluated at 8 $h^{-1}$pc and 50 $h^{-1}$pc. The final column is the age of each model, $t_{\text{now}}$. All parameters are evaluated at this age, including the fluctuation amplitudes.
@@ 3 Idealized voids
An informative starting point for exploring the connection between voids and their background cosmology is an idealized spherical region with a constant density that is, at some early time, slightly less than the average density in the universe. In isolation, this idealized void evolves as if were its own miniature universe with a relatively small matter density and large Hubble parameter [105, 10]. When embedded in a uniform density background universe, we expect the void dynamics to be more complicated as material in the outermost regions of the more rapidly expanding void mixes with the background. Yet the bulk of the idealized void interior evolves as a distinct FLRW world.
We begin this section by considering three approaches to tracking the evolution of idealized voids. The first (section 3.1) comes from linear perturbation theory [84], valid when the amplitude of the relative density contrast is small ($|\delta|\ll 1$). Because the average density contrast never exceeds unity within a void, linear theory offers a useful approximation to the evolution of the matter density there. The second approach (section 3.2) incorporates the Zel’dovich approximation [91], a surprisingly robust quasilinear method [92] for tracking void evolution. Finally, to approach calculating exact solutions, we use the “shell model” in reference [19] to numerically solve the equations of motion for mass elements in an idealized void (section 3.3). These numerical solutions provide an accurate accounting of the mass density even as the outer shells cross into the background universe. Following a brief comparison between these three approaches, we close this section with an assessment how the matter distribution within idealized voids evolve in different background cosmologies.
Throughout, we quantify our results with a void depth parameter [19]
$\varepsilon(r)=1-\frac{\left<\rho\right>_{r}}{\bar{\rho}},$ (3.1)
@@ 3.1 Linear theory
Linear perturbation theory for the growth of cosmic structure [84] describes the evolution of the mass density when relative fluctuations are small, with $|\delta|\ll 1$ (equation (2.10)). Here, we provide the predictions of the theory.
@@
$\varepsilon_{\text{lin}}(z)=\frac{D(z)}{D(z_{\text{init}})}\varepsilon(z_{\text{init}}),$ (3.2)
where redshift $z_{\text{init}}$ is some initial (or reference) redshift, and the denisty growth factors $D$ are defined in equation (2.12). The theory also predicts flow rates for matter within a void. Including both the Hubble flow and the growing-mode solutions from linear theory, the proper radial velocity relative to the center of a void is [84]
$v_{r}={H(z)}r+\frac{2fg}{3{\Omega_{\text{M}}(z)}{H(z)}},$ (3.3)
@@
$=$ $+\frac{{H^{2}(z)}{\Omega_{\text{M}}(z)}r}{2}\varepsilon.$ (3.6)
The void depth $\varepsilon$ in the last equation refers to a tophat underdensity (equation (3.1)) or the amplitude of the mean underdensity within comoving radius $x=r/a$. (Equivalent expressions, as in references [44, 52, 96], are often written in terms of $\Delta(x)$, the mean density contrast within comoving distance $x$.)
An effective Hubble parameter, ${H_{\text{eff}}}$, describes the radial outflow rate of matter in the interior of an idealized void. The linear theory prediction for ${H_{\text{eff}}}$ in units of the background Hubble expansion rate is [84, 106]
$\frac{{H_{\text{eff}}}}{H}$ $=$ $\frac{v_{r}}{r{H(z)}}=1+\frac{2fg}{3{\Omega_{\text{M}}(z)}{H^{2}(z)}r},$ (3.7)
$=$ $1+\frac{f\varepsilon}{3},$ (3.8)
where $r$ and $v_{r}$ are proper distance and speed, respectively. This expression follows from equations (3.3) and (3.6). In section 4 we show that this connection between the outflow rate, the growth factor, and void depth leads to a robust estimator of the background cosmology.